OpenStudy (anonymous):

A 2.44 x 10^3 kg car requires 5.3 kJ of work to move from rest to some final speed. During this time, the car moves 27.4 m. Neglecting friction, find a) the final speed b) the net horizontal force exerted on the car

OpenStudy (anonymous):

for a W = 0.5 m v^2 5300 = 0.5 x 2440 v^2 v = sqrt(4.344) v = 2.08 m/s for b W = F s F = W/s F = 5300/27.4 F =193.43 N

OpenStudy (anonymous):

@lostmermaid happy to helps u :)